OpenStudy (anonymous):

lim as h approaches 0 2x^6+6x^3 over 4x^5+3x^3

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \lim_{h \rightarrow 0}~\frac{2x^6+6x^3 }{4x^5+3x^3} } \]

OpenStudy (solomonzelman):

you meant x not h, right ?

OpenStudy (anonymous):

yes sorry its X

OpenStudy (solomonzelman):

try to factor top and bottom out of x^3, cancel x^3 on top and bottom, and then plug in 0 for x.

OpenStudy (solomonzelman):

\[\huge\color{red}{ \lim_{x \rightarrow 0}~\frac{2x^6+6x^3}{4x^5+3x^3} } \] \[\huge\color{red}{ \lim_{x \rightarrow 0}~\frac{x^3(2x^3+6)}{x^3(4x^2+3)} } \] \[\huge\color{red}{ \lim_{x \rightarrow 0}~\frac{(2x^3+6)}{(4x^2+3)} } \] now plug in zero for x.

OpenStudy (solomonzelman):

What is the answer ?

OpenStudy (anonymous):

2

OpenStudy (solomonzelman):

Yup, that's right.

OpenStudy (anonymous):

thank you so much

OpenStudy (solomonzelman):

You welcome!